# Identification and characterization of distorted gravitational waves by lensing using deep learning

### v1

@@ III.2 Network validation
We perform inference on $10^{3}$ lensed injections generated from our trained model and draw $10^{4}$ samples for each injections from the lensed network. We then reconstruct the reference phase Dax et al. (2023). Figure 2 displays the corresponding probability–probability (P–P) plot constructed directly from the posteriors obtained by the lensed network without applying importance sampling. The diagonal line represents the ideal case in which $x\%$ of the true parameters lie within the $x\%$ credible intervals of the inferred posteriors. The colored curves for individual parameters show excellent agreement with this diagonal, demonstrating that the network produces well-calibrated posteriors. This proper calibration confirms that the sampling procedure functions as intended, a result further supported by the Kolmogorov–Smirnov test $p$-values shown in the legend. Minor deviations for the lensing parameters ($\Delta t$ and $\mu_{\rm rel}$) remain within expected statistical fluctuations ($3\sigma$), indicating robust performance of the model even in the presence of lensing effects.
@@ III.4 Recovery of GWs diffracted by point lenses
Figure 4: P-P plot showing the calibration of the posterior distributions for the lensing parameters $\Delta t$ and $\mu_{\rm rel}$, obtained from $10^{3}$ simulated injections of point-mass–lensed GW signals. Each curve compares the cumulative fraction of true parameter values enclosed within given credible levels against the ideal uniform expectation (black dashed line). The light-gray band indicates the expected $3\sigma$ statistical uncertainty for a perfectly calibrated inference. The analysis uses the geometrical optics approximation to recover signals generated with a point-lens model, demonstrating accurate recovery of both the time delay and relative magnification distributions.
To assess the effectiveness of the geometrical optics approximation in capturing lensing distortions, we perform inference on a lensed GW signal produced by a point-mass lens for which the amplification factor is analytical and depends on the redshifted lens mass $M_{\mathrm{Lz}}$ and impact parameter $y$ Takahashi and Nakamura (2003). The point-mass lens always produces two images, but also includes diffraction effects which are not present in our parametrization of interfering chirps in (3). In practical terms, we use the implementation of the point-mass lens model coded in modwaveforms, [^3] https://github.com/ezquiaga/modwaveforms which is derived on Ref. Ezquiaga et al. (2020). Within the geometric optics limit, the lens mass and impact parameter can be mapped one-to-one with the lensing time delay and relative magnification:
@@
$\mu_{\rm rel}$ $=\left|\frac{y\sqrt{y^{2}+4}-(y^{2}+2)}{y\sqrt{y^{2}+4}+(y^{2}+2)}\right|\,$, (11)
@@
Figure 4 presents the P-P plot assessing the calibration of posterior inferences for the lensing parameters $\Delta t$ and $\mu_{\rm rel}$, based on $10^{3}$ simulated injections of GW signals lensed by a point-mass lens, and $10^{4}$ sample were drawn from the lensed network for each injection. We simulate lensed events by uniformly sampling lens masses from $500M_{\odot}$ to $1500M_{\odot}$. Impact parameters are drawn uniformly from 0 to $1.5$. This choice of parameters ensures that the associated time delay does not exceed $0.1$s. Each curve shows the cumulative fraction of true parameter values contained within the corresponding nominal credible intervals, compared to the ideal uniform expectation indicated by the black dashed line. The light-gray region denotes the expected $3\sigma$ statistical variation for a perfectly calibrated inference under the null hypothesis. While both $\Delta t$ and $\mu_{\rm rel}$ curves exhibit mild deviations beyond the $3\sigma$ envelope, these offsets are small and do not indicate significant systematic biases. Moreover, we also include the corresponding $p$-values from the Kolmogorov-Smirnov test in the legend. This level of agreement demonstrates that, even when using the geometrical optics approximation to recover signals generated with a point-lens model, the inference remains well calibrated and the residual systematics are negligible relative to statistical fluctuations.
@@
Figure 5: The injection corresponds to a point-mass lens with $M_{\mathrm{Lz}}=900M_{\odot}$ and $y=0.3$. The inferred $\Delta t=10.47\pm 0.57\penalty 10000\ \mathrm{ms}$ and $\mu_{\rm rel}=0.66\pm 0.17$ are consistent with the geometrical optics predictions of $\Delta t=11\penalty 10000\ \mathrm{ms}$ and $\mu_{\rm rel}=0.5$, demonstrating the capability of the geometrical optics approximation to accurately reproduce lensing-induced distortions.
@@
To further illustrate the quality of individual recoveries, we focus on a representative example of a lensed signal. Figure 5 presents a subset of the 17-dimensional corner plot and illustrates the recovery of a lensed GW signal using the DINGO-lensing framework for a point-mass lens with $M_{\mathrm{Lz}}=900\penalty 10000\ M_{\odot}$ and $y=0.3$. $10^{5}$ samples have been drawn from the lensed network with utilizing importance sampling. The two-dimensional posterior distributions show tight constraints on the lensing parameters, indicating that the model successfully captures the lens-induced distortions. The inferred $\Delta t=10.47\pm 0.57\penalty 10000\ \mathrm{ms}$ and $\mu_{\rm rel}=0.6\pm 0.17$ are in excellent agreement with the geometrical optics predictions of $\Delta t=11\penalty 10000\ \mathrm{ms}$ and $\mu_{\rm rel}=0.5$, respectively, demonstrating the capability of the geometrical optics approximation to accurately reproduce the lensing effects encoded in the waveform.
@@
We note that this demonstration does not rely on a fully self-consistent point-lensed source model for parameter estimation. Such models, while analytically tractable, neglect waveform modifications arising from finite-source effects and frequency-dependent interference, which become significant outside the geometrical optics regime. Instead, our approach focuses on validating the ability of the DINGO-lensing network to recover the effective observables ($\Delta t$, $\mu_{\rm rel}$) that characterize the lensing distortion within this limit. A direct comparison with full Bayesian recovery using bilby is not performed here, as the inference is expensive and other codes have already implemented this Wright and Hendry (2022); gwm (2023). Our inference test therefore serves as a controlled and efficient validation of the capability of the network to reproduce the predictions of geometrical optics with high fidelity.
@@ III.5 Lensing detectability
Lensing candidates can be identified through different methods. For example, using the posterior samples directly would give an indication of an inconsistency with the nonlensing hypothesis if the lensing parameters deviate away from 0. To quantify this inconsistency, we follow a similar approach to Ref. Ezquiaga et al. (2023) and compute the (Gaussian) distance to $(0,0)$ in the 2D parameter space $(\Delta t,\mu_{\rm rel})$. We define this distance away from the nonlensing hypothesis as $\mathcal{D}_{\mathrm{nonlens}}$, see appendix C for details. Such distance is weighted by the uncertainty of the lensing parameters. Well-constrained posteriors away from 0 would lead to larger distances. Because both lensing parameters are bounded at 0, distances are always larger than 1. We compare the distributions of $\mathcal{D}_{\mathrm{nonlens}}$ and $\mathcal{B}_{\mathrm{lens}}$ in Fig. 6 for 2,561 lensed and 10,858 nonlensed simulations. We find that both are correlated, with lensed simulations displaying a large tail at high values. Noticeably, the distribution of nonlensed simulations has a narrower range of values both in distance and Bayes factors.
We explore in more detail the distribution of Bayes factors in Fig. 7, where we plot the complementary cumulative distribution functions. Lensed simulations display large Bayes factor, with 53% of them being larger than 5. However, for a given lensing candidate, its statistical significance depends on how it compares to a ‘background’, which defines the expectations of the nonlensed hypothesis. In other words, it is necessary to compute the false alarm probability. Similar background studies with DINGO have also been performed in the literature when searching for eccentric signals Gupte et al. (2024).
@@ IV Conclusions
The discovery of lensed GWs requires well-defined detection strategies. DINGO-lensing can contribute to them in two different stages. Firstly, it can be used for the identification of lensed candidates with a set of networks trained on the average power spectral density over a given observing run. These networks should encompass broad mass ranges characterizing different signals duration. The maximum lensing time delay should be adjusted for each network so that there are single, distorted lensed waveforms. Lensing candidates can then be identified through two alternative methods. On one hand, one can compute the distance between the lensing posterior samples and the nonlensing hypothesis, $\Delta t=\mu_{\rm rel}=0$ using only neural posterior estimation. On the other hand, through the use of importance sampling, the Bayes factor between the lensed and nonlensed hypothesis can be compared to pre-computed simulations of nonlensed events. To avoid potential biases, such simulations should describe agnostic models, i.e. uniformly distributed priors in the networks, and astrophysically motivated ones, i.e. matching GW population models with lensing optical depths.
@@
Altogether, DINGO-lensing provides a scalable pathway for rapid lensing analyses in upcoming LVK observing runs. Its efficiency enables large-scale background estimation and systematic population studies that are otherwise computationally prohibitive. Future work will extend this framework to more complex lens models, incorporate waveform systematics, and integrate DINGO-lensing into end-to-end pipelines for real-time identification and characterization of lensed GW events.
@@ Appendix A Nonlensed Network Validation
Figure 9: P-P plot for the nonlensed neural posterior estimation model constructed using $10^{3}$ unlensed injections and no importance sampling. For each injection, a posterior is generated, and the percentile value of the injected parameter is computed. Each colored line corresponds to the cumulative distribution function (CDF) of an individual parameter, with the Kolmogorov–Smirnov (KS) test $p$-values given in the legend.
@@
Figure 10: Posterior distribution of an injection recovered using DINGO (nonlensed network) (green solid curves) and bilby (gray dashed curves), respectively. The two methods agree with each other and both recovered the injected values (solid gray lines; also tabulated in Tab. 1) well.
@@ Appendix C Distance away from the nonlensing hypothesis
For our lens model, cf. Eq. (3), a nonlensed event is characterized by having $\Delta t=\mu_{\rm rel}=0$. Therefore, one can quantify the inconsistency with this hypothesis by measuring the distance of the lensing posterior samples away from this point. We define this quantity as $\mathcal{D}_{\mathrm{nonlens}}$. In order for this distance to be meaningful, it needs to be normalized by the uncertainty of the parameters themselves. Approximating the posterior distributions as Gaussian, the distance in units of $\sigma$ is given by Ezquiaga et al. (2023)
@@
where $\theta_{1}$ and $\theta_{1}$ are the parameters for which the distance is computed, and $C_{1}$ and $C_{2}$ are their covariances. In this case, $\theta_{1}=\{\Delta t,\mu_{\rm rel}\}$ and $\theta_{2}=\{0,0\}$, defining
$\mathcal{D}_{\mathrm{nonlens}}=D(\{\Delta t,\mu_{\rm rel}\},0)\,.$ (14)
Since the prior in $\{\Delta t,\mu_{\rm rel}\}$ is bounded at 0, the lensing distance satisfies $\mathcal{D}_{\mathrm{nonlens}}\gtrsim 1$. It is to be noted that for non-Gaussian posteriors, the Gaussian approximation will typically lead to smaller distances. Therefore, $\mathcal{D}_{\mathrm{nonlens}}$ is a conservative metric to find lensed candidates. The advantage of $\mathcal{D}_{\mathrm{nonlens}}$ over the lensing Bayes factor $\mathcal{B}_{\mathrm{lens}}$ is that it does not require importance sampling and can be obtained directly from neural posterior estimation with DINGO-lensing.

### v2

@@ III.2 Network validation
We perform inference on $10^{3}$ lensed injections generated from our trained model and draw $10^{4}$ samples for each injections from the lensed network. We then reconstruct the reference phase Dax et al. (2023). Figure 2 displays the corresponding probability–probability (P–P) plot constructed directly from the posteriors obtained by the lensed network without applying importance sampling. The diagonal line represents the ideal case in which $x\%$ of the true parameters lie within the $x\%$ credible intervals of the inferred posteriors. The colored curves for individual parameters show excellent agreement with this diagonal, demonstrating that the network produces well-calibrated posteriors. This proper calibration confirms that the sampling procedure functions as intended, a result further supported by the Kolmogorov–Smirnov test $p$-values shown in the legend Seabold et al. (2010). Minor deviations for the lensing parameters ($\Delta t$ and $\mu_{\text{rel}}$) remain within expected statistical fluctuations ($3\sigma$), indicating robust performance of the model even in the presence of lensing effects.
@@ III.4 Recovery of GWs diffracted by point lenses
Figure 4: P-P plot showing the calibration of the posterior distributions for the lensing parameters $\Delta t$ and $\mu_{\text{rel}}$, obtained from $10^{3}$ simulated injections of point-mass–lensed GW signals. Each curve compares the cumulative fraction of true parameter values enclosed within given credible levels against the ideal uniform expectation (black dashed line). The light-gray band indicates the expected $3\sigma$ statistical uncertainty for a perfectly calibrated inference. The analysis uses the geometrical optics approximation to recover signals generated with a point-lens model, demonstrating accurate recovery of both the time delay and relative magnification distributions.
To assess the effectiveness of the geometrical optics approximation in capturing lensing distortions, we perform inference on a lensed GW signal produced by a point-mass lens for which the amplification factor is analytical and depends on the redshifted lens mass $M_{\mathrm{Lz}}$ and impact parameter $y$ Takahashi and Nakamura (2003). The point-mass lens always produces two images, but also includes diffraction effects which are not present in our parametrization of interfering chirps in (3). In practical terms, we use the implementation of the point-mass lens model coded in modwaveforms, which is derived on Ref. Ezquiaga et al. (2020). Within the geometric optics limit, the lens mass and impact parameter can be mapped one-to-one with the lensing time delay and relative magnification:
@@
$\mu_{\text{rel}}$ $=\left|\frac{y\sqrt{y^{2}+4}-(y^{2}+2)}{y\sqrt{y^{2}+4}+(y^{2}+2)}\right|\,$, (11)
@@
Figure 4 presents the P-P plot assessing the calibration of posterior inferences for the lensing parameters $\Delta t$ and $\mu_{\text{rel}}$, based on $10^{3}$ simulated injections of GW signals lensed by a point-mass lens, and $10^{4}$ sample were drawn from the lensed network for each injection. We simulate lensed events by uniformly sampling lens masses from $500M_{\odot}$ to $1500M_{\odot}$. Impact parameters are drawn uniformly from 0 to $1.5$. This choice of parameters ensures that the associated time delay does not exceed $0.1$s, which is the maximum time delay in the training set. Note that because the simulated signal’s duration are always larger than 0.1s, we are testing examples in the wave-optics regime Nakamura (1998); Takahashi and Nakamura (2003). Each curve shows the cumulative fraction of true parameter values contained within the corresponding nominal credible intervals, compared to the ideal uniform expectation indicated by the black dashed line. The light gray region denotes the expected $3\sigma$ statistical variation for a perfectly calibrated inference under the null hypothesis. While both $\Delta t$ and $\mu_{\text{rel}}$ curves exhibit mild deviations beyond the $3\sigma$ envelope, these offsets are small and do not indicate significant systematic biases. Moreover, we also include the corresponding $p$-values from the Kolmogorov-Smirnov test in the legend. This level of agreement demonstrates that, even when using the geometrical optics approximation to recover signals generated with a point-lens model, the inference remains well calibrated and the residual systematics are negligible relative to statistical fluctuations.
@@
Figure 5: The injection corresponds to a point-mass lens with $M_{\mathrm{Lz}}=900M_{\odot}$ and $y=0.3$. The inferred $\Delta t=10.47\pm 0.57\penalty 10000\ \mathrm{ms}$ and $\mu_{\text{rel}}=0.66\pm 0.17$ are consistent with the geometrical optics predictions of $\Delta t=11\penalty 10000\ \mathrm{ms}$ and $\mu_{\text{rel}}=0.5$, demonstrating the capability of the geometrical optics approximation to accurately reproduce lensing-induced distortions.
@@
To further illustrate the quality of individual recoveries, we focus on a representative example of a lensed signal with a duration of $\sim 0.1\rm{s}$. Figure 5 presents a subset of the 17-dimensional corner plot and illustrates the recovery of a lensed GW signal using the DINGO-lensing framework for a point-mass lens with $M_{\mathrm{Lz}}=900\penalty 10000\ M_{\odot}$ and $y=0.3$. $10^{5}$ samples have been drawn from the lensed network with utilizing importance sampling. The two-dimensional posterior distributions show tight constraints on the lensing parameters, indicating that the model successfully captures the lens-induced distortions. The inferred $\Delta t=10.47\pm 0.57\penalty 10000\ \mathrm{ms}$ and $\mu_{\text{rel}}=0.6\pm 0.17$ are in excellent agreement with the geometrical optics predictions of $\Delta t=11\penalty 10000\ \mathrm{ms}$ and $\mu_{\text{rel}}=0.5$, respectively, demonstrating the capability of the geometrical optics approximation to accurately reproduce the lensing effects encoded in the waveform even in the wave-optics regime.
@@
We note that this demonstration does not rely on a fully self-consistent point-lensed source model for parameter estimation. Such models, while analytically tractable, neglect waveform modifications arising from finite-source effects and frequency-dependent interference, which become significant outside the geometrical optics regime. Instead, our approach focuses on validating the ability of the DINGO-lensing network to recover the effective observables ($\Delta t$, $\mu_{\text{rel}}$) that characterize the lensing distortion within this limit. A direct comparison with full Bayesian recovery using bilby is not performed here, as the inference is expensive and other codes have already implemented this Wright and Hendry (2022); gwm (2023). Our inference test therefore serves as a controlled and efficient validation of the capability of the network to reproduce the predictions of geometrical optics with high fidelity.
@@ III.5 Lensing detectability
Lensing candidates can be identified through different methods. For example, using the posterior samples directly would give an indication of an inconsistency with the nonlensing hypothesis if the lensing parameters deviate away from 0. To quantify this inconsistency, we follow a similar approach to Ref. Ezquiaga et al. (2023) and compute the (Gaussian) distance to $(0,0)$ in the 2D parameter space $(\Delta t,\mu_{\text{rel}})$. We define this distance away from the nonlensing hypothesis as $\mathcal{D}_{\mathrm{nonlens}}$, see Appendix C for details. Such distance is weighted by the uncertainty of the lensing parameters. Well-constrained posteriors away from 0 would lead to larger distances. Because both lensing parameters are bounded at 0, distances are always larger than 1. We compare the distributions of $\mathcal{D}_{\mathrm{nonlens}}$ and $\mathcal{B}_{\mathrm{lens}}$ in Fig. 6 for 2,561 lensed and 10,858 nonlensed simulations. We find that both are correlated, with lensed simulations displaying a large tail at high values. Noticeably, the distribution of nonlensed simulations has a narrower range of values both in distance and Bayes factors.
We explore in more detail the distribution of Bayes factors in Fig. 7, where we plot the complementary cumulative distribution functions. Lensed simulations display large Bayes factor, with 53% of them being larger than 5. However, for a given lensing candidate, its statistical significance depends on how it compares to a background, which defines the expectations of the nonlensed hypothesis. In other words, it is necessary to compute the false alarm probability. Similar background studies with DINGO have also been performed in the literature when searching for eccentric signals Gupte et al. (2024).
@@ IV Conclusions
The discovery of lensed GWs requires well-defined detection strategies. DINGO-lensing can contribute to them in two different stages. First, it can be used for the identification of lensed candidates with a set of networks trained on the average power spectral density over a given observing run. These networks should encompass broad mass ranges characterizing different signals duration. The maximum lensing time delay should be adjusted for each network so that there are single, distorted lensed waveforms. Lensing candidates can then be identified through two alternative methods. On one hand, one can compute the distance between the lensing posterior samples and the nonlensing hypothesis, $\Delta t=\mu_{\text{rel}}=0$ using only neural posterior estimation. On the other hand, through the use of importance sampling, the Bayes factor between the lensed and nonlensed hypothesis can be compared to pre-computed simulations of nonlensed events. To avoid potential biases, such simulations should describe agnostic models, i.e. uniformly distributed priors in the networks, and astrophysically motivated ones, i.e. matching GW population models with lensing optical depths.
@@
Altogether, DINGO-lensing provides a scalable pathway for rapid lensing analyses in upcoming LVK observing runs. Its efficiency enables large-scale background estimation and systematic population studies that are otherwise computationally prohibitive. Future work will extend this framework to more complex lens models, incorporate waveform systematics, and integrate DINGO-lensing into end-to-end pipelines for real-time identification and characterization of lensed GW events. Non-gaussianity and nonstationarity in realistic noise can affect both the accuracy of NPE and the detectability of lensing. While the impact on NPE accuracy remains unclear, lensing detectability is likely to be overestimated. A systematic study of NPE performance and lensing detectability under different glitch mitigation techniques, such as BayesWave Cornish and Littenberg (2015); Hourihane et al. (2022) and gwsubtract Davis et al. (2019, 2022), will be carried out in future work.
@@ Appendix A Nonlensed Network Validation
Figure 9: P-P plot for the nonlensed neural posterior estimation model constructed using $10^{3}$ unlensed injections and no importance sampling. For each injection, a posterior is generated, and the percentile value of the injected parameter is computed. Each colored line corresponds to the cumulative distribution function (CDF) of an individual parameter, with the Kolmogorov–Smirnov test $p$-values given in the legend.
@@
Figure 10: Posterior distribution of an injection recovered using DINGO (nonlensed network) (green solid curves) and bilby (gray dashed curves), respectively. The two methods agree with each other and both recovered the injected values (solid gray lines; also tabulated in Table 1) well.
@@ Appendix C Distance away from the nonlensing hypothesis
For our lens model, cf. Eq. (3), a nonlensed event is characterized by having $\Delta t=\mu_{\text{rel}}=0$. Therefore, one can quantify the inconsistency with this hypothesis by measuring the distance of the lensing posterior samples away from this point. We define this quantity as $\mathcal{D}_{\mathrm{nonlens}}$. In order for this distance to be meaningful, it needs to be normalized by the uncertainty of the parameters themselves. Approximating the posterior distributions as Gaussian, the distance in units of $\sigma$ is given by Ezquiaga et al. (2023)
@@
where $\theta_{1}$ and $\theta_{1}$ are the parameters for which the distance is computed, and $C_{1}$ and $C_{2}$ are their covariances. In this case, $\theta_{1}=\{\Delta t,\mu_{\text{rel}}\}$ and $\theta_{2}=\{0,0\}$, defining
$\mathcal{D}_{\mathrm{nonlens}}=D(\{\Delta t,\mu_{\text{rel}}\},0)\,.$ (14)
Since the prior in $\{\Delta t,\mu_{\text{rel}}\}$ is bounded at 0, the lensing distance satisfies $\mathcal{D}_{\mathrm{nonlens}}\gtrsim 1$. It is to be noted that for non-Gaussian posteriors, the Gaussian approximation will typically lead to smaller distances. Therefore, $\mathcal{D}_{\mathrm{nonlens}}$ is a conservative metric to find lensed candidates. The advantage of $\mathcal{D}_{\mathrm{nonlens}}$ over the lensing Bayes factor $\mathcal{B}_{\mathrm{lens}}$ is that it does not require importance sampling and can be obtained directly from neural posterior estimation with DINGO-lensing.